# SN 2024aecx: A double-peaked rapidly evolving Type IIb supernova at 11 Mpc

### v1

@@ 3.4 Spectral evolution
We present the complete spectroscopic evolution of SN 2024aecx in Figure 6. All the spectra have been corrected for host redshift, calibrated with SN photometry, and dereddened for total extinction estimated in Section 3.2. The spectral sequence spans a time period of around 55 days with respect to the explosion epoch. It covers both the shock cooling and secondary peak phases. The initial two spectra at + 1.9 and + 2.0 days show a very blue continuum with some underdeveloped broad features, which are possibly due to hydrogen/silicon and helium. We confirm the presence of H/Si II (6563 Å, 6355 Å) and He (5876 Å) in these early spectra using SYNAPPS modeling, which we discuss later. Below 4000 Å, we observe a prominent emission attributed to Ca II H&K (3934 and 3968 Å). As the ejecta evolves further, we see rapid cooling making the continuum more redder.
Several metallic lines start to form with strong absorptions around +5 day. However, our spectra around +5-6 days have slightly poor SNR than the rest of the spectra owing to the decline in brightness during the shock cooling minima. We find conspicuous features from Fe II and Sc II towards the bluer region. Around 8000–9000 Å, we also observe the development of Ca II triplet. The Na i D line arising due to the host is prominently visible and used to constrain the extinction towards SN (see Section 3.2). Several prominent features identified are marked in the spectra sequence Fig. 6.
@@ 3.5 SYNAPPS models
Owing to uncertainty and ambiguity in some of the identified lines, we attempt to model the early spectrum (+ 1.9 day) and the spectrum around the maximum (+ 22 day) to constrain some of the spectral features. For this purpose, we make use of the open-source SYNAPPS code (Thomas et al., 2011). This is a direct implementation of another code called SYNOW (Parrent et al., 2010), which is a parametrized spectral synthesis code. A spherical symmetry with ejecta being in homologous expansion is assumed in this setup. The photons are emitted from a sharp photosphere, with the optical depth being described as an exponential function of velocity. where is reference velocity for parameterization and is the maximum velocity allowed at the outer edge of the line-forming region (Thomas et al., 2011). For some optical depth, the reference line profile is estimated for a given ion with the remaining lines following Boltzmann statistics (Parrent et al., 2010). SYNAPPS automatically generates synthetic spectra based on a provided set of inputs, including parameters such as ions, blackbody temperature, velocities, and opacities. It continues to iterate over these parameters within pre-defined ranges and compares them with the observed spectra for each iteration to obtain a best-fit model spectrum.
@@
We also model SN 2024aecx spectrum around maximum (), where we find maximum number of elements with well defined line profiles. The best describing model fits are presented in Figure 8. As evident, many of the spectral features are satisfactorily explained with the model fits. The photospheric velocity and temperature obtained at this phase are and , respectively. The region blueward of 5500 Å is mostly dominated by multiple blended features from various metals specifically Fe II, Ti II, and Co II with some contributions from Sc II and Ca II as well. The strong absorption around 5900 Å is well defined by Na I line along with some contribution from He I. There are some other features arising due to He I as seen in the flattened model spectra for He, but due to significant blending from other metal features, it is not trivial to discern the He features in those regions. Similarly, there is some hint of H in the spectra during peak, as we can clearly see H feature around 6500 Åin the model fits. However, we could not trace any other Balmer feature (e.g., H 4861 Å) apart from the weak H feature.
In Fig. 9, we compare SN 2024aecx spectra at similar epochs with other well-studied SNe IIb/Ib indicated in the figure. The comparison spectra (SNe 1993J, 2009jf, 2013df, 2012au, iPTF13bvn, 2022crv) were retrieved from WISeRep (Yaron & Gal-Yam, 2012). The earliest spectrum (+1.9 d) of SN 2024aecx, which belongs to the shock cooling phase, has similarity with SN 1993J and SN 2013df (see left panel of Fig. 9). These are blue and almost featureless. Here, the blue feature is also evident from the early colors as discussed in Section 3.3. However, at a similar epoch, the line formation has already started in SNe 2009jf, 2011fu, iPTF13bvn, and 2022crv. The spectra near the secondary maximum are plotted in the right panel of Fig. 9. By this time, the various spectral lines, such as Fe II, He I, and Sc II can be clearly identified in all events. Notably, the Fe II line near 5000 Å is more prominent in SN 2024aecx than in other events.
@@ 3.5.1 Line velocities
The He I velocity evolved from 12000 to 5000 km s-1 between days 1 to 30 and was always below H velocities. There is a similarity in He I velocity values of SN 2024aecx to other IIb SNe (Liu et al., 2016). Interestingly, in SN 2024aecx the He I line evolved from the beginning; nevertheless, in SN 2013df it started appearing 10 days after the explosion (Morales-Garoffolo et al., 2014). The velocity values of both Fe II (4924 Å, 5018 Å) lines have slower evolution and lie between 7000 to 5000 km s-1. The Ca II 8498 Å have a velocity of 10000 km s-1 around day 30. Both line velocities of Fe II (4924 Å) and Fe II (5018 Å) are consistent. Notably, there is no definite line that can be treated directly for the photosphere; however, Fe II traces it (Dessart & Hillier, 2005). The average of Fe II velocities between day 16 and 22 was considered as the photospheric velocity near maximum and used to calculate the kinetic energy of the explosion in Section 4.
@@ 4 Light curve modeling and explosion parameters
In double-peaked IIb SNe, the propagating shock breaks out the optically thick progenitor envelope after the SN explosion (Falk & Arnett, 1977b). The shock-generated electromagnetic radiation powers the first peak, and a fast decline occurs due to the subsequent cooling of the expanding (extended) envelope (Soderberg et al., 2012). The secondary peak emerges within a few weeks, which is powered by the radioactive decay of 56Ni synthesized in the explosion. The duration of the shock breakout phase (Waxman & Katz, 2017) remains very short (hours) and therefore, usually missed in the observations. The mass of the envelope and the density structure of the outer layers characterize the shock cooling phase. Accordingly, various hydrodynamical and analytical models have been proposed to understand the evolution of IIb SNe progenitors and the explosion. It includes helium stars with a thin hydrogen envelope (see Shigeyama et al., 1994; Woosley et al., 1994; Blinnikov et al., 1998; Rabinak & Waxman, 2011; Nakar & Piro, 2014; Bersten et al., 2012b; Katz et al., 2012; Bersten et al., 2018b).
@@
The extended progenitors and low-mass envelopes of the double-peaked SN 1993J and SN 2011dh were explained with theoretical modeling (Nakar & Piro, 2014). These calculations were successively improved in later studies by incorporating a two-zone model of variable power-law density for better modeling of the early shock cooling features (see Piro et al., 2021). Further, Sapir & Waxman (2017) refined the shock cooling model considering a complicated polytropic density profile, which was missing in previous models. Recently, Morag et al. (2023) proposed another model for similar studies by combining previous model solutions of Sapir et al. (2013) and Sapir & Waxman (2017).
@@
Both models have some limitations and approximations. Namely, the temperature validity is only for 0.7 eV 8120 K for both models. Also, the valid envelope mass range is 0.1 and 1 10 for Sapir & Waxman (2017) and Morag et al. (2023), respectively. Here, it is emphasized that the Sapir & Waxman (2017) model is preferred over Morag et al. (2023) as the previous is more suitable for SNe IIb due to a better acceptable range (but see Farah et al., 2025). Therefore, we use the Sapir et al. (2013) model in our further analysis, which is implemented in the open-source package shock-cooling-curve (Venkatraman & Jacobson-Galán, 2024). The code follows a Markov Chain Monte Carlo (MCMC) routine. Considering that the temperature criteria (see Fig. 4, top panel) are satisfied up to less than 5 days of the shock cooling phase, and hence, the model was fitted to this phase only.
@@
Both polytropic indices n = 3 and n = 3/2, which signify the radiative and convective envelopes, were executed. The MCMC routine was operated to achieve the best fit major parameters such as envelope mass (Men), progenitor radius (R), and shock velocity scale (vsh). It also estimates the explosion time (t0). The best fitted parameters were achieved after numerous runs (2000), and the resulting model is plotted in Fig. 11. The estimated parameter R, Men, vsh and t0 values are: 169.19 R☉, 0.24 M☉, 2.35 and – 0.55 days (for polytropic index n = 3); and 200.35 R☉, 0.03 M☉, 1.55 and – 1.00 days (for polytropic index n = 3/2), respectively. These are listed in Table 4. The corner plots resulted from the fitting are shown in Appendix References (Fig. 13 and 14).
@@
We used bolometric light curves computed in Section 3.3 to estimate the explosion parameters of SN 2024aecx. The conceptualizations given by (Arnett, 1982) and (Valenti et al., 2008) were used to fit the second peak, which is mainly believed to be powered by the radioactive decay of 56Ni. The models are based on several assumptions that the ejecta are spherically symmetric and homologous, the opacity () is constant, and 56Ni is centrally located and unmixed (Arnett, 1982; Valenti et al., 2008; Cano, 2013). The parameters and are the nickel mass and diffusion time-scale, respectively, and were kept as free variables in the fitting. The ejecta kinetic energy and are represented by
@@
Here, is a constant of integration (Valenti et al., 2008), vph is photospheric velocity, and is the speed of light. The optical opacity is a constant (Chugai, 2000; Toy et al., 2016; Taddia et al., 2018a). The fits were performed with the transient fitting model implemented in Redback (Sarin et al., 2024). In the fitting, the ejecta mass , fraction of nickel fNi, representative initial ejecta velocity vej, , the gamma-ray opacity were kept as free parameters. The estimated values of the explosion parameters are: = 0.15 0.06 M☉, = 0.70 M☉ and = 0.16 1051 erg. was estimated using the above equation, and the photospheric velocity was taken as 6000 km s-1 (see Section 3.4).
@@ 5 Discussion and Summary
The early phase multi-band observations of double-peaked IIb SNe with a good cadence are limited in the literature. Existing large survey facilities are restricted to a single or double band on a particular night. The temporal evolution of the shock breakout, followed by the shock cooling, is rapid, and the duration depends on the progenitor structure. Therefore, observations in the early phase after the explosion are challenging. Because of their peculiar observational properties, SNe IIb are regarded as gap-bridging events between IIP and Ib/c. Hence, additional study of such SNe can play a crucial role in understanding stellar evolution and supernova physics. We investigated simultaneous multi-band imaging of SN 2024aecx, including the shock cooling phase. Our analysis infers that the duration of the shock cooling phase is longer in bluer bands, but the rise time to the secondary peak (after the explosion) is shorter in those bands. Also, the second peak is broader in redder bands. It is further noted that the shock cooling phase of SN 2024aecx is shorter than several Type IIb events (SNe 1993J, 2011fu and 2013df).
@@
Recently, Ayala et al. (2025) performed a detailed statistical study of early light curves of IIb SNe based on a sample from the ATLAS survey and well-studied literature data. They categorize these events into two groups: EE (an early flux excess SNe followed by a secondary peak) and non-EE (without early flux excess SNe). SN 2024aecx falls to their EE group (cf. double peak). In Fig. 12, the trise (rise time to secondary peak after the explosion) and m15 of their sample and the values estimated for SN 2024aecx are displayed (also see Table 3). There is a clear trend that events with faster post-peak decline rates attain a brighter secondary peak earlier (Ayala et al., 2025). It is worth noting that SN 2024aecx has very fast evolution among the sample, which is also evident from Figure 5. Nevertheless, such an inference can be confirmed with an additional data sample in the future. Further, the shock cooling duration of SN 2024aecx is significantly less than the mean duration (8.85 days) as estimated in Ayala et al. (2025). With a peak absolute magnitude (second peak) of Mg = –17.94 mag, SN 2024aecx is one of the brightest events in SNe IIb sample, but it does not follow the trend that brighter SNe exhibit longer rise times; rather, it peaked earlier. Similarly, the color evolution of SN 2024aecx is unique, and various colors display a trend of blue-red and red-blue-red between 0–8 days and 8–40 days. Such a feature is likely because of the interplay between the shock cooling and the cooling of the expanding ejecta. Unlike SNe IIn, no precursor activity appeared in SN during the monitoring of NGC 3521 for more than a year.
@@
The spectroscopic properties of SN 2024aecx have similarities with typical IIb SNe, but the hydrogen line features are weak and could only be seen up to about 1 month after the explosion. The He I line evolved from the beginning against SN 2013df, where it started appearing 10 days past-explosion. To discern several of the spectral features, we performed a crude modeling of SN 2024aecx using SYNAPPS. A significant number of lines were identified and corroborated, justifying the usual spectral properties of IIb events. However, few spectral lines remain evasive. With this modeling, not much could be said about such spectral features, and a thorough analysis of the ejecta composition and evolution could be carried out using more sophisticated spectral modeling codes (e.g., CMFGEN, TARDIS, etc.), which is beyond the scope of this work.
@@
The shock cooling phase multi-band data ( 5 days) of SN 2024aecx were modeled with Sapir & Waxman (2017) shock cooling analytical model. The radiative (n = 3) and convective (n = 3/2) polytropic indices were used to infer the progenitor properties (see Section 4). The estimated best-fit envelope mass has a range of 0.03–0.24 , indicating a moderately stripped progenitor. This is similar to SN 1993J progenitor, where Me was estimated less than 0.4 (Woosley et al., 1994). Such a smaller envelope mass also corroborates the weak hydrogen appearance in SN 2024aecx spectra. Furthermore, the progenitor radius is 169 R☉ and 200 R☉ for n=3/2 and 3. This range has a similarity with progenitor radii of SN 2016gkg (Tartaglia et al., 2017b; Piro et al., 2017). The overall observational properties and analytical modeling parameters of SN 2024aecx indicates that its progenitor was possibly less extended than SN 1993J (600 R☉ Maund et al., 2004) and SN 2011fu (450 R☉ Kumar et al., 2013; Morales-Garoffolo et al., 2015) and SN 2013df (550 R☉ Van Dyk et al., 2014). However, Morales-Garoffolo et al. (2014) estimated a smaller radius for SN 2013df (64–169 R☉). The intermediate extended progenitor of SN 204aecx was also favored considering a shorter shock cooling time of SN 2024aecx (Section 3). Nevertheless, a more robust hydrodynamical modeling can provide a more clear picture of SN 2024aecx progenitor system.
@@
The modeling of pseudo-bolometric light curve implies an ejecta mass of 0.70 M☉, an explosion energy 0.16 1051 erg, and the synthesized 56Ni 0.15 M☉. The production of 56Ni in SE-SNe is correlated with the explosion properties and the structure of the progenitor core (see Suwa et al., 2019, and references therein). Higher 56Ni mass signifies a more luminous event, which supports the brighter peak of SN 2024aecx. However, the 56Ni evaluation using Arnett’s rule may overestimate the 56Ni value in SE-SNe as Arnett’s model is more appropriate for Type Ia SNe (Dessart et al., 2015, 2016; Khatami & Kasen, 2019). Recent sample-based studies suggest 56Ni production range of 0.060 – 0.124 for Type IIb SNe (Anderson, 2019; Prentice et al., 2019; Afsariardchi et al., 2021; Rodríguez et al., 2023). The estimated 56Ni value for SN 2024aecx is consistent with Anderson (2019) and also with the derived 56Ni values for SN 2011fu (Kumar et al., 2013) and SN 2013df (Morales-Garoffolo et al., 2014). The light curve shape of SE-SNe is also influenced by the mixing of 56Ni in the ejecta (Ensman & Woosley, 1988; Woosley & Eastman, 1997; Shigeyama et al., 1990; Cano et al., 2014; Dessart et al., 2015, 2016). The 56Ni mixing may arise due to various reasons, e.g., large-scale asymmetric explosion jets and asymmetry of the propagating shock, etc. (Maund et al., 2009; Dessart et al., 2012). The early color evolution can suggest the degree of mixing. The monotonic and non-monotonic color evolution hints at strong and weak mixing, respectively (Yoon et al., 2019; Moriya et al., 2020). SN 2024aecx early colors, which display a non-monotonic evolution, may be attributed to a weak to moderate 56Ni mixing. Further, as the ejecta mass of SN 2024aecx is smaller, it will result shorter diffusion time of the radiation (Rodríguez et al., 2023) and so faster light curve evolution.

### v2

@@ 3.4 Spectral evolution
We present the complete spectroscopic evolution of SN 2024aecx in Figure 6. All the spectra have been corrected for host redshift, calibrated with SN photometry, and dereddened for total extinction estimated in Section 3.2. The spectral sequence spans a time period of around 55 days with respect to the explosion epoch. It covers both the shock cooling and secondary peak phases. The initial two spectra at + 1.9 and + 2.0 days show a very blue continuum with some underdeveloped broad features, which are possibly due to hydrogen/silicon and helium. We confirm the presence of H/Si II (6563 Å, 6355 Å) and He (5876 Å) in these early spectra using SYNAPPS modeling, which we discuss later. Below 4000 Å, we observe a prominent emission attributed to Ca II H&K (3934 and 3968 Å). As the ejecta evolves further, we see rapid cooling, making the continuum redder.
Several metallic lines start to form with strong absorptions around +5 day. However, our spectra around +5-6 days have slightly poorer SNR than the rest of the spectra owing to the decline in brightness during the shock cooling minima. We find conspicuous features from Fe II and Sc II towards the bluer region. Around 8000–9000 Å, we also observe the development of Ca II triplet. The Na i D line arising due to the host is prominently visible and used to constrain the extinction towards SN (see Section 3.2). Several prominent features identified are marked in the spectra sequence Fig. 6.
@@ 3.5 SYNAPPS models
Owing to uncertainty and ambiguity in some of the identified lines, we attempt to model the early spectrum (+ 1.9 day) and the spectrum around the maximum (+ 22 day) to constrain some of the spectral features. For this purpose, we make use of the open-source SYNAPPS code (R. C. Thomas et al., 2011). This is a direct implementation of another code called SYNOW (J. Parrent et al., 2010), which is a parametrized spectral synthesis code. A spherical symmetry with ejecta being in homologous expansion is assumed in this setup. The photons are emitted from a sharp photosphere, with the optical depth being described as an exponential function of velocity. where is reference velocity for parameterization and is the maximum velocity allowed at the outer edge of the line-forming region (R. C. Thomas et al., 2011). For some optical depth, the reference line profile is estimated for a given ion with the remaining lines following Boltzmann statistics (J. Parrent et al., 2010). SYNAPPS automatically generates synthetic spectra based on a provided set of inputs, including parameters such as ions, blackbody temperature, velocities, and opacities. It continues to iterate over these parameters within pre-defined ranges and compares them with the observed spectra for each iteration to obtain a best-fit model spectrum.
@@
We also model the SN 2024aecx spectrum around maximum (), where we find the maximum number of elements with well-defined line profiles. The best describing model fits are presented in Figure 8. As is evident, many of the spectral features are satisfactorily explained with the model fits. The photospheric velocity and temperature obtained at this phase are and , respectively. The region blueward of 5500 Å is mostly dominated by multiple blended features from various metals, specifically Fe II, Ti II, and Co II with some contributions from Sc II and Ca II as well. The strong absorption around 5900 Å is well defined by Na I line along with some contribution from He I. There are some other features arising due to He I as seen in the flattened model spectra for He, but due to significant blending from other metal features, it is not trivial to discern the He features in those regions. Similarly, there is some hint of H in the spectra during peak, as we can clearly see the H feature around 6500 Å in the model fits. However, we could not trace any other Balmer feature (e.g., H 4861 Å) apart from the weak H feature.
In Fig. 9, we compare SN 2024aecx spectra at similar epochs with other well-studied SNe Ib, Ic, and IIb indicated in the figure. The comparison spectra (SNe 1993J, 1994I, 2007gr, 2009jf, 2013df, 2012au, iPTF13bvn, 2022crv) were retrieved from the respective studies (see Section 3.1 for references) and from WISeRep (O. Yaron & A. Gal-Yam, 2012). The earliest spectrum (+1.9 d) of SN 2024aecx, which belongs to the shock cooling phase, has similarity with SN 1993J and SN 2013df (see left panel of Fig. 9). These are blue and almost featureless. Here, the blue feature is also evident from the early colors as discussed in Section 3.3. However, at a similar epoch, line formation has already begun in SNe 2007gr, 2009jf, 2011fu, iPTF13bvn, and 2022crv. The spectra near the secondary maximum are plotted in the right panel of Fig. 9. By this time, the various spectral lines, such as Fe II, He I, and Sc II can be clearly identified in all events. Notably, the Fe II line near 5000 Å is more prominent in SN 2024aecx than in other events.
@@ 3.5.1 Line velocities
The He I velocity evolved from 12000 to 5000 km s-1 between days 1 to 30 and was always below H velocities. There is a similarity in He I velocity values of SN 2024aecx to other IIb SNe (Y.-Q. Liu et al., 2016). Interestingly, in SN 2024aecx the He I line evolved from the beginning; nevertheless, in SN 2013df it started appearing 10 days after the explosion (A. Morales-Garoffolo et al., 2014). The velocity values of both Fe II (4924 Å, 5018 Å) lines have slower evolution and lie between 7000 to 5000 km s-1. The Ca II 8498 Å have a velocity of 10000 km s-1 around day 30. Both line velocities of Fe II (4924 Å) and Fe II (5018 Å) are consistent. Notably, there is no definite line that can be treated directly for the photosphere; however, Fe II traces it (L. Dessart & D. J. Hillier, 2005). The average of Fe II velocities between day 16 and 22 was considered as the photospheric velocity near maximum and used to calculate the kinetic energy of the explosion in Section 4.
@@ 4 Light curve modeling and explosion parameters
In double-peaked IIb SNe, the propagating shock breaks out the optically thick progenitor envelope after the SN explosion (S. W. Falk & W. D. Arnett, 1977b). The shock-generated electromagnetic radiation powers the first peak, and a fast decline occurs due to the subsequent cooling of the expanding (extended) envelope (A. M. Soderberg et al., 2012). The secondary peak emerges within a few weeks, which is powered by the radioactive decay of 56Ni synthesized in the explosion. The duration of the shock breakout phase (E. Waxman & B. Katz, 2017) remains very short ( hours) and is therefore usually missed in observations. The mass of the envelope and the density structure of the outer layers characterize the shock cooling phase. Accordingly, various hydrodynamical and analytical models have been proposed to understand the evolution of IIb SNe progenitors and the explosion. It includes helium stars with a thin hydrogen envelope (see T. Shigeyama et al., 1994; S. E. Woosley et al., 1994; S. I. Blinnikov et al., 1998; I. Rabinak & E. Waxman, 2011; E. Nakar & A. L. Piro, 2014; M. C. Bersten et al., 2012b; B. Katz et al., 2012; M. C. Bersten et al., 2018b).
@@
The extended progenitors and low-mass envelopes of the double-peaked SN 1993J and SN 2011dh were explained with theoretical modeling (E. Nakar & A. L. Piro, 2014). These calculations were successively improved in later studies by incorporating a two-zone model of variable power-law density for better modeling of the early shock cooling features (see A. L. Piro et al., 2021). Further, N. Sapir & E. Waxman (2017) refined the shock cooling model, considering a complicated polytropic density profile, which was missing in previous models. Recently, J. Morag et al. (2023) proposed another model for similar studies by combining previous model solutions of N. Sapir et al. (2013) and N. Sapir & E. Waxman (2017).
@@
Both models have some limitations and approximations. Namely, the temperature validity is only for 0.7 eV 8120 K for both models. Also, the valid envelope mass range is 0.1 and 1 10 for N. Sapir & E. Waxman (2017) and J. Morag et al. (2023), respectively. Here, it is emphasized that the N. Sapir & E. Waxman (2017) model is preferred over J. Morag et al. (2023) as the previous is more suitable for SNe IIb due to a better acceptable range (but see J. R. Farah et al., 2025). Therefore, we use the N. Sapir et al. (2013) model in our further analysis, which is implemented in the open-source package shock-cooling-curve (P. Venkatraman & W. Jacobson-Galán, 2024). The code follows a Markov Chain Monte Carlo (MCMC) routine. Considering that the temperature criteria (see Fig. 5, top panel) are satisfied up to less than 5 days of the shock cooling phase, and hence, the model was fitted to this phase only.
@@
Both polytropic indices n = 3 and n = 3/2, which signify the radiative and convective envelopes, were executed. The MCMC routine was operated to achieve the best fit major parameters such as envelope mass (Men), progenitor radius (R), and shock velocity scale (vsh). It also estimates the explosion time (t0). The best-fitted parameters were achieved after numerous runs (2000), and the resulting model is plotted in Fig. 11. The estimated parameter R, Men, vsh and t0 values are: 169.19 R☉, 0.24 M☉, 2.35 109 cm s-1 and – 0.55 days (for polytropic index n = 3); and 200.35 R☉, 0.03 M☉, 1.55 109 cm s-1 and – 1.00 days (for polytropic index n = 3/2), respectively. These are listed in Table 4. The corner plots resulting from the fitting are shown in Appendix References (Fig. 13 and 14).
@@
We used bolometric light curves computed in Section 3.3 to estimate the explosion parameters of SN 2024aecx. The conceptualizations given by (W. D. Arnett, 1982) and (S. Valenti et al., 2008) were used to fit the second peak, which is mainly believed to be powered by the radioactive decay of 56Ni. The models are based on several assumptions that the ejecta are spherically symmetric and homologous, the opacity () is constant, and 56Ni is centrally located and unmixed (W. D. Arnett, 1982; S. Valenti et al., 2008; Z. Cano, 2013). The parameters and are the nickel mass and diffusion time-scale, respectively, and were kept as free variables in the fitting. The ejecta kinetic energy and are represented by
@@
Here, is a constant of integration (S. Valenti et al., 2008), vph is photospheric velocity, and is the speed of light. The optical opacity is a constant (N. N. Chugai, 2000; V. L. Toy et al., 2016; F. Taddia et al., 2018a). The fits were performed with the transient fitting model implemented in Redback (N. Sarin et al., 2024). In the fitting, the ejecta mass , fraction of nickel fNi, representative initial ejecta velocity vej, , the gamma-ray opacity were kept as free parameters. The estimated values of the explosion parameters are: = 0.15 0.06 M☉, = 0.70 M☉ and = 0.16 1051 erg. was estimated using the above equation, and the photospheric velocity was taken as 6000 km s-1 (see Section 3.4).
@@ 5 Discussion and Summary
The early phase multi-band observations of double-peaked IIb SNe with a good cadence are limited in the literature. Existing large survey facilities are restricted to a single or double band on a particular night. The temporal evolution of the shock breakout, followed by the shock cooling, is rapid, and the duration depends on the progenitor structure. Therefore, observations in the early phase after the explosion are challenging. Because of their peculiar observational properties, SNe IIb are regarded as gap-bridging events between IIP and Ib/c. Hence, additional study of such SNe can play a crucial role in understanding stellar evolution and supernova physics. We investigated simultaneous multi-band imaging of SN 2024aecx, including the shock cooling phase. Our analysis infers that the duration of the shock cooling phase is longer in bluer bands, but the rise time to the secondary peak (after the explosion) is shorter in those bands. Also, the second peak is broader in redder bands. It is further noted that the shock cooling phase of SN 2024aecx is shorter than several Type IIb events (SNe 1993J, 2011fu, and 2013df).
@@
Recently, B. Ayala et al. (2025) performed a detailed statistical study of early light curves of IIb SNe based on a sample from the ATLAS survey and well-studied literature data. They categorize these events into two groups: EE (an early flux excess SNe followed by a secondary peak) and non-EE (without early flux excess SNe). SN 2024aecx falls to their EE group (cf. double peak). In Fig. 12, the trise (rise time to secondary peak after the explosion) and m15 of their sample and the values estimated for SN 2024aecx are displayed (also see Table 3). There is a clear trend that events with faster post-peak decline rates attain a brighter secondary peak earlier (B. Ayala et al., 2025). It is worth noting that SN 2024aecx has very fast evolution among the sample, which is also evident from Figure 3. Nevertheless, such an inference can be confirmed with an additional data sample in the future. Further, the shock cooling duration of SN 2024aecx (see Table 3) is significantly less than the mean duration (8.85 days) as estimated in B. Ayala et al. (2025). With a peak absolute magnitude (second peak) of Mg = –17.94 mag, SN 2024aecx is one of the brightest events in the SNe IIb sample, but it does not follow the trend that brighter SNe exhibit longer rise times; rather, it peaked earlier. Similarly, the color evolution of SN 2024aecx is unique, and various colors display a trend of blue-red and red-blue-red between 0 – 8 days and 8 – 40 days. Such a feature is likely because of the interplay between the shock cooling and the cooling of the expanding ejecta. During the monitoring of NGC 3521 for more than a year with Mephisto, the forced photometry at the SN location does not exhibit any reliable flux enhancement.
@@
The spectroscopic properties of SN 2024aecx have similarities with typical IIb SNe, but the hydrogen line features are weak and could only be seen up to about 1 month after the explosion. To discern several of the spectral features, we performed a crude spectral modeling of SN 2024aecx using SYNAPPS. A significant number of lines were identified and corroborated, justifying the usual spectral properties of IIb events (see Section 3.5). However, a few spectral lines remain evasive. The presence of the hydrogen line in the early spectra of SN 2024aecx is an important clue to differentiate it from Type Ib events. However, the possible contamination arising due to the blending of Si II line near 6500 Å can not be ignored.
Another important line He I 5876 Å, evolved from the beginning in SN 2024aecx against SN 2013df, where it started appearing 10 days post-explosion. Nevertheless, the trough near this line could be blended with Na I (5890 Å, 5986 Å). In Type Ic SNe, the identification of He I 5876 Å line is debatable (A. Clocchiatti et al., 1996; T. Matheson et al., 2001; A. Elmhamdi et al., 2006; D. N. Sauer et al., 2006; L. Dessart et al., 2015a; D. Milisavljevic et al., 2015). Y.-Q. Liu et al. (2016) analyzed a large optical spectroscopic data of Ib and Ic SNe. In their study, no convincing signs of either He I 6678 Å or 7065 Å were noticed in the mean spectra of Ic SNe. They suggested no obvious He I lines in Ic SNe (see also M. Modjaz et al., 2016). The He I 10830 Å and 20581 Å lines in the NIR region are indeed useful for an affirmation between the helium-rich and helium-poor events (M. Shahbandeh et al., 2022). However, we can not comment on this without NIR spectra of SN 2024aecx. It is worth mentioning that there are limitations with the spectral modeling presented in this paper. A thorough analysis of the ejecta composition and evolution could be carried out using more sophisticated spectral modeling codes (e.g., CMFGEN, TARDIS, etc.), which is beyond the scope of this work.
@@
The shock cooling phase multi-band data ( 5 days) of SN 2024aecx were modeled with N. Sapir & E. Waxman (2017) shock cooling analytical model. The radiative (n = 3) and convective (n = 3/2) polytropic indices were used to infer the progenitor properties (see Section 4). The estimated best-fit envelope mass has a range of 0.03 – 0.24 , indicating a moderately stripped progenitor. This is similar to SN 1993J progenitor, where Me was estimated less than 0.4 (S. E. Woosley et al., 1994). Such a smaller envelope mass also corroborates the weak hydrogen appearance in SN 2024aecx spectra. Furthermore, the progenitor radius is 169 R☉ and 200 R☉ for n = 3 and 3/2, respectively. This range has a similarity with progenitor radii of SN 2016gkg (L. Tartaglia et al., 2017; A. L. Piro et al., 2017). The overall observational properties and analytical modeling parameters of SN 2024aecx indicates that its progenitor was possibly less extended than SN 1993J ( 600 R☉ J. R. Maund et al., 2004), SN 2011fu (450 R☉ B. Kumar et al., 2013; A. Morales-Garoffolo et al., 2015) and SN 2013df (550 R☉ S. D. Van Dyk et al., 2014). However, A. Morales-Garoffolo et al. (2014) estimated a smaller radius for SN 2013df (64 – 169 R☉). The intermediate extended progenitor of SN 204aecx was also favored, considering a shorter shock cooling time of SN 2024aecx (see Section 3). Nevertheless, a more robust hydrodynamical modeling can provide a more clear picture of the SN 2024aecx progenitor system.
@@
The modeling of pseudo-bolometric light curve implies an ejecta mass of 0.70 M☉, an explosion energy 0.16 1051 erg, and the synthesized 56Ni 0.15 M☉. The production of 56Ni in SE-SNe is correlated with the explosion properties and the structure of the progenitor core (see Y. Suwa et al., 2019, and references therein). Higher 56Ni mass signifies a more luminous event, which supports the brighter peak of SN 2024aecx. However, the 56Ni evaluation using Arnett’s rule may overestimate the 56Ni value in SE-SNe as Arnett’s model is more appropriate for Type Ia SNe (L. Dessart et al., 2015b, 2016; D. K. Khatami & D. N. Kasen, 2019). Recent sample-based studies suggest 56Ni production range of 0.060 – 0.124 for Type IIb SNe (J. P. Anderson, 2019; S. J. Prentice et al., 2019; N. Afsariardchi et al., 2021; Ó. Rodríguez et al., 2023).
The estimated 56Ni value for SN 2024aecx is consistent with J. P. Anderson (2019) and also with the derived 56Ni values for SN 2011fu (B. Kumar et al., 2013) and SN 2013df (A. Morales-Garoffolo et al., 2014). The light curve shape of SE-SNe is also influenced by the mixing of 56Ni in the ejecta (L. M. Ensman & S. E. Woosley, 1988; S. E. Woosley & R. G. Eastman, 1997; T. Shigeyama et al., 1990; Z. Cano et al., 2014; L. Dessart et al., 2015b, 2016). The 56Ni mixing may arise due to various reasons, e.g., large-scale asymmetric explosion jets and asymmetry of the propagating shock, etc. (J. R. Maund et al., 2009; L. Dessart et al., 2012). The early color evolution can suggest the degree of mixing. The monotonic and non-monotonic color evolution hints at strong and weak mixing, respectively (S.-C. Yoon et al., 2019; T. J. Moriya et al., 2020). SN 2024aecx early colors, which display a non-monotonic evolution, may be attributed to a weak to moderate 56Ni mixing. Further, as the ejecta mass of SN 2024aecx is smaller, it will result shorter diffusion time of the radiation (Ó. Rodríguez et al., 2023) and so faster light curve evolution.